# Provably Robust Score-Based Diffusion Posterior Sampling for Plug-and-Play Image Reconstruction

### v1

@@ Abstract
This work develops an algorithmic framework for employing score-based diffusion models as an expressive data prior in nonlinear inverse problems with general forward models. Motivated by the plug-and-play framework in the imaging community, we introduce a diffusion plug-and-play method (DPnP) that alternatively calls two samplers, a proximal consistency sampler based solely on the likelihood function of the forward model, and a denoising diffusion sampler based solely on the score functions of the image prior. The key insight is that denoising under white Gaussian noise can be solved rigorously via both stochastic (i.e., DDPM-type) and deterministic (i.e., DDIM-type) samplers using the same set of score functions trained for generation. We establish both asymptotic and non-asymptotic performance guarantees of DPnP, and provide numerical experiments to illustrate its promise in solving both linear and nonlinear image reconstruction tasks. To the best of our knowledge, DPnP is the first provably-robust posterior sampling method for nonlinear inverse problems using unconditional diffusion priors.
@@ 2.1 The forward process and score functions
To facilitate understanding, it will be convenient to formulate the diffusion process in continuous time. To distinguish from the discrete-time setting, we use capitalized letters like ${X}$ to denote the continuous-time diffusion process, and $\tau$ to denote continuous time parameter. The continuous-time forward diffusion follows the Ornstein-Uhlenbeck (OU) process, defined by the Stochastic Differential Equation (SDE) (Song et al., 2021b, ):
@@ 2.2 The reverse process and sampling
    $\mathrm{d}{M}^{\mathsf{rev}}_{\tau^{\mathsf{rev}}}=\left({M}^{\mathsf{rev}}_{\tau^{\mathsf{rev}}}+\nabla\log p_{{M}_{\tau^{\mathsf{rev}}}}(\tau^{\mathsf{rev}},{M}^{\mathsf{rev}}_{\tau^{\mathsf{rev}}})\right)\mathrm{d}\tau,\quad\tau\in[0,\tau_{\infty}],\quad{M}^{\mathsf{rev}}_{\tau_{\infty}}\sim p_{{M}_{\tau_{\infty}}}.$ (13)
@@ 3.3 Our algorithm: diffusion plug-and-play
Discussions of DPnP.
Some comments about the proposed DPnP method are in order.
@@ 5 Numerical experiments
We provide preliminary numerical evidence to corroborate the promise of DPnP in solving both linear and nonlinear image reconstruction tasks. We leave more comprehensive experiments to future work.
Input DPS DPnP-DDPM DPnP-DDIM Ground truth
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Samples of different algorithms for super-resolution, where DPnP-DDPM and DPnP-DDIM recover fine details of the image more faithfully than the state-of-the-art DPS algorithm (Chung et al., 2023a, ).
@@
We discuss a heuristic strategy to choose the annealing schedule $\eta_{k}$ in DPnP (Algorithm 3). As seen in the theoretical analysis (Theorem 2), if we set all the $\eta_{k}\equiv\eta$ for some constant $\eta>0$, then DPnP converges to a distribution $\pi_{\eta}$, which can be regarded as a version of the posterior distribution $p^{\star}(\cdot|y)$ distorted by an order of $O(\eta)$. The smaller $\eta$ is, the more accurate the final distribution will be. On the other hand, it was also seen that in many cases, the spectral gap is $\Omega(\eta)$, hence the convergence time is $O(\frac{1}{\eta})$. Therefore, smaller $\eta$ would make it take longer to converge. [^5] Strictly speaking, while the number of iterations required to converge increases as $\eta$ gets smaller, the computational complexity of PCS and DDS per iteration will decrease. However, in experiments, we observed that the latter effect is not strong enough to compensate for the increase in overall complexity caused by the former.
@@
where $K_{0}<K$ is the length of the initial phase, which can be chosen as, e.g. $K_{0}=K/5$.
Input DPS DPnP-DDPM DPnP-DDIM Ground truth
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Samples of different algorithms for phase retrieval, where DPnP-DDPM and DPnP-DDIM recover fine details of the image more faithfully than the state-of-the-art DPS algorithm (Chung et al., 2023a, ).
We denote DPnP with the subroutines DDS-DDPM and DDS-DDIM as DPnP-DDPM and DPnP-DDIM respectively. We compare DPnP with the state-of-the-art DPS algorithm (Chung et al., 2023a, ) for image super-resolution and phase retrieval.
  • Image super-resolution. For super-resolution, the forward model is the bicubic downsampling operator, which is a linear operator. The samples generated by different algorithms are shown in Table 1.
  • Phase retrieval. For phase retrieval, the forward model is as follows. Fourier transform is performed to the image with a coded mask, and only the magnitude of the Fourier transform is taken as the measurement (Shechtman et al.,, 2015). Due to the nonlinearity of the operation of taking magnitude, the forward model is nonlinear. The samples generated by different algorithms are shown in Table 2.
It can be seen from these results that, DPnP is capable of solving both linear and nonlinear problems, and, in comparison with DPS, performs better in recovering fine and crisper details.

### v2

@@ Abstract
This work develops an algorithmic framework for employing score-based diffusion models as an expressive data prior in nonlinear inverse problems with general forward models. Motivated by the plug-and-play framework in the imaging community, we introduce a diffusion plug-and-play method (DPnP) that alternatively calls two samplers, a proximal consistency sampler based solely on the likelihood function of the forward model, and a denoising diffusion sampler based solely on the score functions of the image prior. The key insight is that denoising under white Gaussian noise can be solved rigorously via both stochastic (i.e., DDPM-type) and deterministic (i.e., DDIM-type) samplers using the same set of score functions trained for generation. We establish both asymptotic and non-asymptotic performance guarantees of DPnP, and provide numerical experiments to illustrate its promise in solving both linear and nonlinear image reconstruction tasks. To the best of our knowledge, DPnP is the first provably-robust posterior sampling method for nonlinear inverse problems using unconditional diffusion priors. Code is available at https://github.com/x1xu/diffusion-plug-and-play.
@@ 2.1 The forward process and score functions
To facilitate understanding, it will be convenient to formulate the diffusion process in continuous time. To distinguish from the discrete-time setting, we use capitalized letters like ${X}$ to denote the continuous-time diffusion process, and $\tau$ to denote continuous time parameter. The continuous-time forward diffusion follows the Ornstein-Uhlenbeck (OU) process [^1] In the literature, other processes such as Variance-Exploding SDE (VE-SDE) are also used. Our theory also applies to these processes with straightforward modifications., defined by the Stochastic Differential Equation (SDE) (Song et al., 2021b, ):
@@ 2.2 The reverse process and sampling
    $\mathrm{d}{M}^{\mathsf{rev}}_{\tau^{\mathsf{rev}}}=\left(-f({M}^{\mathsf{rev}}_{\tau^{\mathsf{rev}}})+\frac{\beta}{2}\nabla\log p_{{M}_{\tau^{\mathsf{rev}}}}(\tau^{\mathsf{rev}},{M}^{\mathsf{rev}}_{\tau^{\mathsf{rev}}})\right)\mathrm{d}\tau,\quad\tau\in[0,\tau_{\infty}],\quad{M}^{\mathsf{rev}}_{\tau_{\infty}}\sim p_{{M}_{\tau_{\infty}}}.$ (13)
@@ 3.3 Our algorithm: diffusion plug-and-play
Discussions of DPnP.
Some comments about the proposed DPnP method are in order.
  • We note that the proximal consistency sampler PCS in fact admits a simple form in the special case where the forward model $\mathcal{A}$ is linear, i.e. $\mathcal{A}(x)=Ax$ for some matrix $A\in\mathbb{R}^{m\times d}$ (assume $A$ is of full row rank), and the measurement noise $\xi\sim\mathcal{N}(0,\Sigma)$ is Gaussian. In this situation, the log-likelihood function becomes
    $\mathcal{L}(x;y)=-\frac{1}{2}(y-Ax)^{\top}\Sigma^{-1}(y-Ax)+\text{constant}.$
    Then the target distribution of PCS can be rewritten as a Gaussian distribution with proper mean and covariance, namely,
    $\exp\left(\mathcal{L}(\cdot;{y})-\frac{1}{2\eta_{k}^{2}}\|\cdot-\widehat{x}_{k}\|^{2}\right)\propto\exp\left(-\frac{1}{2}(\cdot-\tilde{x}_{k})^{\top}\tilde{\Sigma}_{k}^{-1}(\cdot-\tilde{x}_{k})\right)=\mathcal{N}(\widetilde{x}_{k},\widetilde{\Sigma}_{k}),$
    where
    $\widetilde{x}_{k}=\left(A^{\top}\Sigma^{-1}A+\frac{1}{\eta_{k}^{2}}I_{d}\right)^{-1}\left(A^{\top}\Sigma^{-1}y+\frac{1}{\eta_{k}^{2}}\widehat{x}_{k}\right),\qquad\widetilde{\Sigma}_{k}=\left(A^{\top}\Sigma^{-1}A+\frac{1}{\eta_{k}^{2}}I_{d}\right)^{-1}.$
    Therefore, the proximal consistency sampler PCS can be implemented directly by
    $\widehat{x}_{k+\frac{1}{2}}=\textsf{PCS}(\widehat{x}_{k},y,\mathcal{L},\eta_{k})=\widetilde{x}_{k}+\widetilde{\Sigma}_{k}^{1/2}{w}_{k},\quad{w}_{k}\sim\mathcal{N}(0,I_{d}).$
@@ 5 Numerical experiments
We provide numerical evidence to corroborate the promise of DPnP in solving both linear and nonlinear image reconstruction tasks. We denote DPnP with the subroutines DDS-DDPM and DDS-DDS-DDIM as DPnP-DDPM and DPnP-DDS-DDIM respectively.
5.1 Inverse problems
We consider the following linear and nonlinear inverse problems in our experiments.
Phase retrieval.
We consider phase retrieval with a coded mask, which is a classical inverse problem (Candes et al.,, 2015). For a $256\times 256$ image $x$ (for each color channel) in our experiments, we first generate a random mask $M\in\mathbb{R}^{256\times 256}$ (which is shared across color channels), then apply Fourier transform $\mathcal{F}$ to $M\odot x$, where $\odot$ denotes the Hadamard (entrywise) product, and finally preserve only the magnitudes of the Fourier transform. Formally, the forward measurement operator is $\mathcal{A}(x)=\operatorname{mag}(\mathcal{F}(M\odot x))$, where $\operatorname{mag}(\cdot)$ computes the entrywise magnitude of a matrix with complex entries. The measurement noise is again set to be white Gaussian, with variance $0.2$.
Quantized sensing.
Quantized sensing refers to the task of reconstructing an image from its low-bit quantized version. Here, the forward measurement operator is a one-bit per channel, dithered quantization operator. More precisely, it applies entrywise the following stochastic function $Q$ with dithering level $\theta>0$:
$Q(\mathsf{pixel})=\begin{cases}1,&\text{with probability }\frac{\mathrm{e}^{\mathsf{pixel}/\theta}}{1+\mathrm{e}^{\mathsf{pixel}/\theta}}\\ -1,&\text{with probability }\frac{1}{1+\mathrm{e}^{\mathsf{pixel}/\theta}},\end{cases}$
where $\mathsf{pixel}\in[-1,1]$ is the value of each pixel in each channel. The measurements in quantized sensing are therefore one-bit-per-channel images. The dithering level $\theta$ is set to $0.4$ in our experiments.
Super resolution.
The forward model for super-resolution is the bicubic downsampling operator (Keys,, 1981), which is a linear operator (in fact, a block Hankel matrix). We use a downsampling ratio of $4$ in all our experiments. The measurement noise is set to be white Gaussian, with variance $0.2$.
5.2 Experimental setups
We compare DPnP with the state-of-the-art DPS algorithm (Chung et al., 2023a, ) and LGD-MC algorithm (Song et al., 2023b, ) on the FFHQ validation dataset (Karras et al.,, 2019) and the ImageNet validation dataset (Russakovsky et al.,, 2015). We use the same pre-trained score functions as in Chung et al., 2023a , [^6] https://github.com/DPS2022/diffusion-posterior-sampling and all images are normalized to fit into the range $[-1,1]$.
Figure 1: Samples of different algorithms for phase retrieval, quantized sensing, and super resolution, where DPnP generate images of higher quality and recover fine details of the image more faithfully than the state-of-the-art DPS (Chung et al., 2023a, ) and LGD-MC (Song et al., 2023b, ) algorithms.
Input DPS LGD-MC DPnP-DDPM DPnP-DDIM Ground truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) phase retrieval
Input DPS LGD-MC DPnP-DDPM DPnP-DDIM Ground truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(b) quantized sensing
Input DPS LGD-MC DPnP-DDPM DPnP-DDIM Ground truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c) super resolution
@@ 5.2 Experimental setups
For DPnP, we use a heuristic strategy to choose the annealing schedule $\eta_{k}$ in DPnP (Algorithm 3). As seen in the theoretical analysis (Theorem 2), if we set all the $\eta_{k}\equiv\eta$ for some constant $\eta>0$, then DPnP converges to a distribution $\pi_{\eta}$, which can be regarded as a version of the posterior distribution $p^{\star}(\cdot|y)$ distorted by an order of $O(\eta)$. The smaller $\eta$ is, the more accurate the final distribution will be. On the other hand, it was also seen that in many cases, the spectral gap is $\Omega(\eta)$, hence the convergence time is $O(\frac{1}{\eta})$. Therefore, smaller $\eta$ would make it take longer to converge. [^7] Strictly speaking, while the number of iterations required to converge increases as $\eta$ gets smaller, the computational complexity of PCS and DDS per iteration will decrease. However, in experiments, we observed that the latter effect is not strong enough to compensate for the increase in overall complexity caused by the former.
@@
where $K_{0}<K$ is the length of the initial phase, which can be chosen as, e.g., $K_{0}=K/5$. For all the numerical experiments, we set $\eta_{0}=0.4$, $\eta_{N}=0.15$, $K_{0}=4$, $K=20$. The annealing schedule $\{\eta_{k}\}$ of DPnP is fixed across all tasks, while DPS and LGD-MC are fine-tuned with reasonable effort for best performance (see Appendix D.1 for the choice of hyperparameters). All experiments are run on a single Nvidia L40 GPU.
Initialization.
In Algorithm 3, the initial guess $\widehat{x}_{0}$ is set to be a properly scaled Gaussian random vector. Notwithstanding, from Theorem 2 it can be infered that using a heuristic posterior sampler as the initializer could decrease $\chi^{2}(p_{\widehat{x}_{1}}\,\|\,\pi_{\eta})$, hence potentially improve the convergence speed of DPnP. By using existing algorithms like DPS or LGD-MC as initializers, DPnP can improve upon the results of existing algorithms towards the correct posterior distribution efficiently and provably. In our experiments, we find it helpful to initialize DPnP with LGD-MC, which accelerates the algorithm significantly.
5.3 Results
Visual results.
The samples generated by different algorithms are shown in Figure 1. It can be seen from these results that, DPnP is capable of solving both linear and nonlinear problems, and, in comparison with state-of-the-art algorithms, performs better in recovering fine and crisper details.
Performance metric.
We report the performance metric of DPnP in terms of LPIPS and PSNR — which are two of the most relevant metrics for inverse problems — on the FFHQ and ImageNet datasets in Table 1 and Table 2, respectively. Since DPnP-DDIM has similar performance with DPnP-DDPM but admits much faster implementation, only DPnP-DDIM is evaluated. We refer to comparisons to other metrics such as FID and SSIM to Appendix D.2. It can be seen that DPnP-DDIM performs strongly on both datasets, albeit taking about 1.5x more computation time.
Table 1: Evaluation of solving inverse problems on FFHQ $256\times 256$ validation dataset (1k samples).
Super-resolution Phase retrieval Quantized sensing Time
(4x, linear) (nonlinear) (nonlinear) per sample
Algorithm LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$
DPnP-DDIM (ours) $\mathbf{0.301}$ $\mathbf{24.2}$ $\mathbf{0.376}$ $\mathbf{22.4}$ $\mathbf{0.293}$ $\mathbf{24.2}$ $\sim 90$s
DPS (Chung et al., 2023a, ) $0.331$ $23.1$ $0.490$ $17.4$ $0.367$ $21.7$ $\sim 60$s
LGD-MC ($n=5$) (Song et al., 2023b, ) $0.318$ $23.9$ $0.522$ $16.4$ $0.317$ $23.9$ $\sim 60$s
Table 2: Evaluation of solving inverse problems on ImageNet $256\times 256$ validation dataset (1k samples).
Super-resolution Phase retrieval Quantized sensing Time
(4x, linear) (nonlinear) (nonlinear) per sample
Algorithm LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$
DPnP-DDIM (ours) $\mathbf{0.416}$ $\mathbf{21.6}$ $\mathbf{0.562}$ $\mathbf{13.4}$ $\mathbf{0.363}$ $\mathbf{23.0}$ $\sim 240$s
DPS (Chung et al., 2023a, ) $0.473$ $20.2$ $0.677$ $\mathbf{13.4}$ $0.542$ $18.7$ $\sim 150$s
LGD-MC ($n=5$) (Song et al., 2023b, ) $\mathbf{0.416}$ $20.9$ $0.592$ $12.8$ $0.384$ $22.3$ $\sim 150$s
@@ C.6 Proof of Lemma 9
Appendix D Additional experiment details
D.1 Hyperparameters of competing methods
We made our best effort to fine-tune the other algorithms within a reasonable amount of time for each task. We list the parameters, following the notation in the original papers (Chung et al., 2023a, ; Song et al., 2023b, ), as follows.
  • Phase retrieval: For DPS, the stepsize is set to $0.8$. For LGD-MC, the MC sampling variance $r_{t}=0.05$, the loss coefficient $\lambda=10^{-3}$, and the learning rate is set to $400.0$.
  • Quantized sensing: For DPS, the stepsize is set to $100.0$. For LGD-MC, the MC sampling variance $r_{t}=0.05$, the loss coefficient $\lambda=2\times 10^{-5}$, and the learning rate is set to $500.0$.
  • Super-resolution: For DPS, the learning rate is set to $0.6$. For LGD-MC, the MC sampling variance $r_{t}=0.05$, the loss coefficient $\lambda=10^{-3}$, and the learning rate is set to $60.0$.
D.2 Additional performance metrics
Computation time in terms of Neural Function Estimations (NFEs).
In additional to the clock time statistics, we also measure the computational cost per sample of different algorithms in terms of the number of Neural Function Estimations, i.e., the number of calls to score functions. The results are in Table 3. Note that the NFEs for DPnP depends on the initialization, the annealing schedule (and the number of timesteps for DPnP-DDIM). We provide typical numbers of NFEs with the choice of parameters used in this paper.
Table 3: Number of NFEs for different algorithms.
Algorithm DPnP-DDIM DPnP-DDPM DPS LGD-MC
NFEs $\sim 1500$ $\sim 3000$ $1000$ $1000$
Frechet Inception Distance (FID) and Structural Similarity Index Measure (SSIM).
We also compare the FID and SSIM of different algorithms across different tasks. The results are shown in Table 4 and Table 5. It should be pointed out that FID is arguably not very relevant to measure the quality of solving inverse problems, as accurately solving inverse problems means that the generated distribution is close to the conditional, i.e., posterior distribution of the image, while FID only measure the closeness to the unconditional, i.e., prior distribution of the image.
Table 4: FID and SSIM of solving inverse problems on FFHQ $256\times 256$ validation dataset (1k samples).
Super-resolution Phase retrieval Quantized sensing
(4x, linear) (nonlinear) (nonlinear)
Algorithm FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$
DPnP-DDIM (ours) $\mathbf{36.3}$ $\mathbf{0.668}$ $\mathbf{46.5}$ $\mathbf{0.631}$ $\mathbf{37.3}$ $\mathbf{0.712}$
DPS (Chung et al., 2023a, ) $38.6$ $0.636$ $52.0$ $0.494$ $42.1$ $0.601$
LGD-MC ($n=5$) (Song et al., 2023b, ) $36.8$ $0.651$ $82.3$ $0.414$ $40.3$ $0.639$
Table 5: FID and SSIM of solving inverse problems on ImageNet $256\times 256$ validation dataset (1k samples).
Super-resolution Phase retrieval Quantized sensing
(4x, linear) (nonlinear) (nonlinear)
Algorithm FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$
DPnP-DDIM (ours) $47.5$ $\mathbf{0.510}$ $\mathbf{73.5}$ $0.289$ $\mathbf{43.2}$ $\mathbf{0.623}$
DPS (Chung et al., 2023a, ) $61.4$ $0.496$ $92.7$ $\mathbf{0.318}$ $82.4$ $0.459$
LGD-MC ($n=5$) (Song et al., 2023b, ) $\mathbf{46.2}$ $0.503$ $89.8$ $0.234$ $46.8$ $0.563$